OpenStudy (anonymous):

ill medal && fan you! just answer my question! :o find the value of x.. 2x+(x+15)= ?

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

(?-3)/15 = x

OpenStudy (anonymous):

2x+(x+15)=? 2x+x+15=? 3x+15=? /3 x+5=? thats as far as you can go really, without the = value you cant truely solve it.

OpenStudy (anonymous):

2x+x=-15 3x=-15 x=(-15)/3 x=-5

OpenStudy (anonymous):

if you say the ? is 0

OpenStudy (anonymous):

there was no equal sign actually.. sorry!

OpenStudy (anonymous):

3x+15